Evaluate the following integrals:
ii) ∫ ln(x+1)/(√x+1)dx
1
Expert's answer
2019-04-29T10:21:35-0400

Solution. We use the integration formula in parts


"\int \frac {ln(x+1)} {\sqrt{x+1}}dx=
\begin{vmatrix}
 u=ln(x+1) & dv=\frac {dx} {\sqrt{x+1}} \\
 du=\frac {dx}{x+1} & v=2\sqrt{x+1}
\end{vmatrix}="

"=2\sqrt{x+1}*ln(x+1)-\int \frac{2\sqrt{x+1}}{x+1}dx="

"=2\sqrt{x+1}*ln(x+1)-\int \frac{2} {\sqrt{x+1}}dx="


"=2\sqrt{x+1}*ln(x+1)-4\sqrt{x+1}+C"

where C is constant.

Answer.


"2\sqrt{x+1}*ln(x+1)-4\sqrt{x+1}+C"

where C is constant.